# supplement.pdf

### v1

@@ 3.4 Material and Lighting Optimization
In this section, we focus on detailing our optimization process and the optimization function. Given the object mesh, our method adaptively create the floor and four walls in the scene of the differentiable renderer, and adaptively places initial area lights on each wall and above the object. The intensity of these lights is optimizable. We set the light bounce count to 2 to enable global illumination. Our optimization process consists of two stages. In the first stage, we keep the lighting fixed and optimize only the parameters of the procedural materials. In the second stage, we jointly optimize the lighting and the procedural materials.
We denote the rendering function of the differentiable renderer PSDR-CUDA as $R(\cdot)$, and represent the parameters of all assigned procedural materials for the mesh as $\theta$. To obtain the UV maps, we use “Smart UV projection” of Blender for each material part. Given the rendered images $R(\theta)_{i}$ and the reference images $I_{i}$, where $i$ refers to the camera view index, we calculate the loss (Eq. 2) between them to backpropagate the gradient from the pixel level to the computational graph of procedural materials, thereby optimizing the parameters $\theta$. We use Adam [21] as the optimizer, and the total optimization function is,
$\mathcal{L}_{\text{total}}=\lambda_{1}\mathcal{L}_{\text{resized}}+\lambda_{2}\mathcal{L}_{\mathrm{\text{pixel}}}+\lambda_{3}\mathcal{L}_{\mathrm{\text{stat}}}+\lambda_{4}\mathcal{L}_{\mathrm{\text{gram}}},\vspace{-0.1cm}$ (2)
where $\lambda_{1-4}$ are the weights, $\mathcal{L}_{\text{resized}}$ is the $\mathcal{L}_{1}$ difference between the reference image and rendered image (downsized to $1/8$), $\mathcal{L}_{\text{pixel}}$ is the average $\mathcal{L}_{1}$ difference between the reference image and rendered image with each material part mask, $\mathcal{L}_{\text{stat}}$ is the mean absolute difference on the statistics (mean $\mu$ and variance $\sigma^{2}$) of the masked pixels of each material part between the two images, and $\mathcal{L}_{\text{gram}}$ is the average $\mathcal{L}_{1}$ difference on the Gram matrix texture descriptor [11] $T_{g}$ for each part. The four sub-optimization objectives are defined as follows:
@@
$\mathcal{L}_{\text{stat}}=\frac{1}{\sum_{i=1}^{n}m_{i}}\sum_{i=1}^{n}\sum_{j=1}^{m_{i}}|\mu(M_{ij}*R(\theta)_{i})\!-\!\mu(M_{ij}*I_{i})|\\ +|\sigma^{2}(M_{ij}*R(\theta)_{i})\!-\!\sigma^{2}(M_{ij}*I_{i})|,$ (5)
$\mathcal{L}_{\text{gram}}=\frac{1}{\sum_{i=1}^{n}m_{i}}\sum_{i=1}^{n}\sum_{j=1}^{m_{i}}\\ \mathcal{L}_{1}(T_{g}(M_{ij}*R(\theta)_{i}),T_{g}(M_{ij}*I_{i})),$ (6)
where $i$ represents the camera view index, $n$ represents the total number of camera views, $j$ refers to the material part index, $m_{i}$ refers to the total number of material parts under camera view $i$, $R(\theta)_{i}^{\prime}$ and $I_{i}^{\prime}$ are the downsized images, and $M_{ij}$ is the mask of the material part $j$.
@@ 4 Experiments
4.1 Experiment Setup
Datasets. We experimented our method on the 3D-FUTURE [10], 3DCoMPaT [26] and Objaverse [8] dataset. For detailed datasets description, please refer to our supplementary.
Baselines. We compare our method with two SOTA text-driven texture generation methods, TEXTure [33] and Text2Tex [6]. These two approaches utilize the Stable Diffusion [34] to generate 2D images, which are then projected onto the mesh. They generate complete textures through multiple iterations using an in-painting approach. However, they only produce diffuse textures with baked lighting. As a result, they cannot support relighting, so in the relighting experiment of Sec. 4.3, we only present the results of our method. Please refer to our supplementary video for intuitive qualitative results.
@@ 4.2 Comparison to Baselines
In this section, we show the qualitative and quantitative comparison results with baselines. Similar to TEXTure [33], we conducted a user study with the same configuration to quantitatively evaluate the quality of text-driven texture generation. Specifically, we prepare 10 testing examples (shapes and text descriptions) for each method. We ask 30 participants to rate the overall quality and the text fidelity, on a scale from 1 to 5. The results are shown in Table. 1, indicating that our method achieves the best score among all the methods. In Fig. 6, we qualitatively compare the texturing results of different methods from both front and back viewpoints. As shown in Fig. 6, both TEXTure and Text2Tex have some artifacts, as both methods require projecting the generated 2D images onto the mesh based on the inpainting approach, resulting in missing textures in some obscured areas. Moreover, due to the lack of multi-view datasets training, the Stable Diffusion model utilized by these two methods can result in inconsistencies across different views, thus causing the Janus (multi-faced) problem. In contrast, ours can generate textures that maintain consistency across multiple views and produce photorealistic texturing results. The more texturing results are shown in Fig. 7.
Figure 7: More Texturing Results. For each sample, we present four different viewpoints. We recommend zooming in to view the texturing results.
@@ 4.3 Relighting Results
Figure 9: Different Point Light Positions. The top row corresponds to point light positioned northeast, and the bottom row corresponds to southeast from three viewpoints.
@@ 4.4 Ablation Studies
Material agent for material classification. Through material agent, we can achieve object-aware material classification with the order of possibility. Moreover, the number of predicted class labels is unconstrained. Here, we compare against a simple baseline. This baseline calculates the VGG loss for all materials in the library using Eq. 1 (with $\alpha=1$), then selects the 10 materials with the lowest loss and uses a voting way to determine the most frequently occurring material type as the predicted type. We present the comparison results in Fig. 10, where the upper part shows part-level understanding by agent and the lower part shows the comparison of material classification for different material parts.
Final aligned masks. For a material part, accurately obtaining its mask on the reference image is crucial for guiding material agent and calculating the optimization function (Eq. 2). As shown in Fig. 11, since the reference image is not well aligned with the mesh geometry, the rendered mask cannot be directly applied to the reference image. Thus, we propose to obtain the final aligned mask by intersecting Matcher [27] mask and rendered mask.
@@
Figure 11: Final Aligned Mask. We show the three material parts of a chair. We can’t directly apply the rendered mask or the Matcher mask to the reference image.
@@ 5 Conclusion
We present a new method to generate high-fidelity materials for 3D shapes given textual prompts. Unlike previous texture generation methods that generate RGB textures containing baked lighting, our method can generate diverse texture maps that enable physical-based rendering and relighting. Additionally, our approach is based on procedural material modeling, which allows us to inherit many advantages of procedural materials. Currently, our method requires pre-segemented material parts which can be improved by integrating with the shape analysis methods [13, 22, 39] in the future.

### v2

@@ 3.5 Material and Lighting Optimization
In this section, we detail our optimization process and the optimization function. Given the object mesh, our method adaptively creates the floor and four walls in the scene of the differentiable renderer, and adaptively places initial area lights on each wall and above the object. The intensity of these lights is included in the optimization. We set the light bounce count to 2 to enable global illumination. Our optimization process consists of two stages. In the first stage, we keep the lighting fixed and optimize only the parameters of the procedural materials. In the second stage, we jointly optimize the lighting and the procedural materials.
We denote the rendering function of the differentiable renderer PSDR-CUDA as $R(\cdot)$, and represent the parameters of all assigned procedural materials for the mesh as $\theta$. To obtain the $U$-$V$ maps, we use ‘Smart UV projection’ in Blender for each material part. Given the rendered images $R(\theta)_{i}$ and the reference images $I_{i}$, where $i$ indexes the camera view, we calculate the loss in Eq. 2 between them to backpropagate the gradient from the pixel level to the computational graph of procedural materials, thereby optimizing the parameters $\theta$. We use Adam [42] as the optimizer. The total optimization objective is:
$\mathcal{L}_{\text{total}}=\lambda_{1}\mathcal{L}_{\text{resized}}+\lambda_{2}\mathcal{L}_{\mathrm{\text{pixel}}}+\lambda_{3}\mathcal{L}_{\mathrm{\text{stat}}}+\lambda_{4}\mathcal{L}_{\mathrm{\text{gram}}},$ (2)
where $\lambda_{1-4}$ are weights, $\mathcal{L}_{\text{resized}}$ is the $\mathcal{L}_{1}$ difference between the reference image and rendered image (downsized to $1/8$), $\mathcal{L}_{\text{pixel}}$ is the average $\mathcal{L}_{1}$ difference between the reference image and rendered image with each material part mask, $\mathcal{L}_{\text{stat}}$ is the mean absolute difference of the statistics (mean $\mu$ and variance $\sigma^{2}$) of the masked pixels of each material part between the two images, and $\mathcal{L}_{\text{gram}}$ is the average $\mathcal{L}_{1}$ difference of the Gram matrix texture descriptor [43] $T_{g}$ for each part. These four sub-objectives are defined as follows:
@@
{multline}
L_stat=1∑i=1nmi ∑_i=1^n∑_j=1^m_i
μ(M_ij*R(θ)_i)-μ(M_ij*I_i)—+
σ^2 (M_ij*R(θ)_i)-σ^2 (M_ij*I_i)—, {multline} L_gram=1∑i=1nmi ∑_i=1^n∑_j=1^m_i
L_1(T_g(M_ij*R(θ)_i),T_g(M_ij*I_i)), where $i$ represents the camera view index, $n$ represents the total number of camera views, $j$ refers to the material part index, $m_{i}$ refers to the total number of material parts under camera view $i$, $R(\theta)_{i}^{\prime}$ and $I_{i}^{\prime}$ are downsized images, and $M_{ij}$ is the mask for material part $j$.
@@ 4 Experiments
4.1 Experimental Setup
Datasets. We performed experiments using the 3D-FUTURE [30], 3DCoMPaT [31] and Objaverse [33] datasets. For detailed datasets description, please refer to the appendix.
Baselines. We compared our method with two state of the art text-driven texture generation methods, TEXTure [2] and Text2Tex [1].
These two approaches utilize Stable Diffusion [3] to generate 2D images, which are then projected onto the mesh. They generate complete textures through multiple iterations using an in-painting approach. However, they only produce diffuse textures with baked lighting. As a result, they cannot support relighting, so in the relighting experiment of Sec. 4.3, we only present the results of our method. Please refer to our supplementary video for intuitive qualitative results.
@@ 4.2 Comparison to Baselines
This section provides qualitative and quantitative comparisons with the baselines. Following TEXTure [2], we conducted a user study with the same configuration to quantitatively evaluate the quality of text-driven texture generation. Specifically, we prepared 10 test examples (shapes and text descriptions) for each method. We asked 30 participants to rate the overall quality and fidelity to the text, on a scale from 1 to 5. The results are shown in Table 1, indicating that our method achieves the best scores among all methods. In Fig. 6, we qualitatively compare the texturing results of different methods from both front and back viewpoints. Both TEXTure and Text2Tex produce some artifacts, as both methods require projecting the generated 2D images onto the mesh based on the inpainting approach, resulting in missing textures in some obscured areas. Moreover, due to the lack of multi-view datasets training, the Stable Diffusion model utilized by these two methods can result in inconsistencies across different views, thus causing the Janus (multi-faced) problem. In contrast, our method can generate textures that maintain consistency across multiple views and produce photorealistic texturing results. More texturing results are shown in Fig. 7.
Figure 7: More Texturing Results. For each sample, we present four different viewpoints.
@@ 4.3 Relighting Results
Figure 9: Different Point Light Positions. The top row corresponds to a point light positioned northeast, and the bottom row corresponds to southeast, each shown from three viewpoints.
@@ 4.4 Ablation Studies
Material agent for material classification. Through the material agent, we can achieve object-aware material classification with the order of likelihood. Moreover, the number of predicted class labels is unconstrained. Here, we compare against a simple baseline. This baseline calculates the VGG loss for all materials in the library using Eq. 1 (with $\alpha=1$), then selects the 10 materials with the lowest loss and uses voting to determine the most frequently occurring material type as the predicted type. We present the comparison results in Fig. 10, where the upper part shows part-level understanding by the agent and the lower part compares the material classification for different material parts.
@@
Final aligned masks. For a material part, accurately obtaining its mask on the reference image is crucial for guiding the material agent and calculating the optimization function (Eq. 2). As shown in Fig. 11, since the reference image is not well aligned with the mesh geometry, the rendered mask cannot be directly applied to the reference image. Thus, we obtain the final aligned mask by intersecting the Matcher [7] mask and rendered mask.
Figure 11: Final Aligned Mask. We show the three material parts of a chair. We cannot directly apply the rendered mask or the Matcher mask to the reference image.
5 Limitation
Our approach also has some limitations. Since it optimizes the parameters of individual nodes within the procedural material graph, it is fundamentally constrained by the materials available in the procedural material library. As a result, it is currently incapable of generating complex textures, such as letters or characters. Nevertheless, if such textures already exist within the library, our method can still support their generation. A potential future direction could involve a two-stage generation process, where complex textures like characters are synthesized and overlaid on top of procedural materials, further expanding the expressiveness of our approach.
@@ 6 Conclusion
We present a new method to generate high-fidelity materials for 3D shapes given textual prompts. Unlike previous texture generation methods that generate RGB textures containing baked lighting, our method can generate diverse texture maps that enable physically-based rendering and relighting. Additionally, our approach is based on procedural material modeling, which allows us to inherit many advantages of procedural materials. Currently, our method requires pre-segmented material parts which could be improved by integrating it with shape analysis methods [35, 36, 37] in the future.
Appendix
This appendix provides further details and results omitted from the main paper. Specifically, we provide additional implementation details in Sec. A and experimental details in Sec. B. In Sec. C, we present more comparisons involving our method. Please refer to our supplementary video for more vivid method explanations and qualitative results.
Appendix A Implementation Details
A.1 Camera Views Selection
Given two elevation angles and the distance to the object, we use spherical sampling to uniformly sample $n$ camera coordinates for each elevation. We then construct the extrinsic camera parameters using the “Look At” approach. We calculate the number of pixels for each material part under each viewpoint through the rendered mask $M_{R}$. We select a sufficient number of cameras from these $2\times n$ cameras to ensure that the pixel count for each part exceeds a threshold (set to 500). To deal with the slight differences between reference images from different views, we select only one reference image for each material part.
Specifically, the front view of the larger elevation will first be selected, because, using this view, the reference image generated by Stable Diffusion [3] will be more realistic and most material parts will exceed the pixel count threshold.
For material parts with a pixel count below the threshold under this front view, including those unseen parts, we then adaptively select $k$ cameras based on their pixel counts. To balance the goals of minimizing the number of selected cameras and maximizing the pixel count for each material part, we adopt three steps to select the cameras for the remaining material parts. The first step is to select the “unavoidable” cameras. For each material part, we compare the maximum pixel count under the $12n-$ 1 viewpoints with the threshold. If the maximum value is less than the threshold, we select the camera corresponding to the maximum pixel count to ensure sufficient visibility for this material part. Through this step, we select $n_{1}$ viewpoints. Next, we select cameras for the remaining material parts. The second step is to select the minimal subset of cameras that satisfy the threshold condition. We mark the cameras for each material part where the pixel count exceeds the threshold, then select the minimal subset of marked cameras to ensure that each material part meets the threshold condition. Through this step, we select $n_{2}$ viewpoints. The third step involves finding the intersection of the initially selected front view, the $n_{1}$ viewpoints selected in the first step, and the $n_{2}$ viewpoints selected in the second step, and using the resulting intersection as the final set of chosen cameras. Through this step, we can obtain the final $k$ cameras. Note that during optimization, each material part will only use the viewpoint selected for it.
Figure A1: Prompt template.
A.2 Smoothing Matcher Masks
Matcher [7] can segment anything by using an in-context example without training. As shown in Fig. 2 of the main paper, we use the rendering of the object textured by white material under specific ambient lighting and the mask $M_{R}$ of a material part on this rendered image as the in-context example. Then, the Matcher will output the corresponding mask $M_{I}$ for the reference image.
We experimented with different ambient lighting to illuminate the objects for the in-context example. In the end, we choose the indoor uniform white light for the ambient lighting, because the segmentation of Matcher will be more accurate.
The masks output by Matcher typically contain noise. Therefore, we subsequently smooth the masks through three image processing steps. We first apply median filtering to attempt to connect regions and remove some outliers. In the second step, we use the morphological erosion operation on the binary images (masks) to remove some outlier areas. Finally, for each Matcher mask $M_{I}$, we count the number of pixels $\{P_{1},\dots,P_{i},\dots\}$ in each connected region $\{R_{1},\dots,R_{i},\dots\}$, where $P_{i}$ is the number of pixels of region $R_{i}$. We calculate the smallest number of pixels $P_{r}$ for all connected regions in the rendered mask $M_{R}$. We then set the connected regions with a pixel count smaller than $\beta P_{r}$ ($\beta=0.5$) to False (i.e., Masked) for every region in $\{R_{1},\dots,R_{i},\dots\}$.
A.3 Material Agent Prompt Design
Prompt image. We choose the reference image generated by Stable Diffusion [3] to construct the prompt image rather than the rendered image of the object textured by white material under specific ambient lighting, because the former carries more information, leading to more accurate part-level understanding by MLLMs. To support part-level understanding, we tested four object labeling schemes: (i) obtaining contours of the material part through mask $M_{ij}$, then outlining the contours with red lines; (ii) building upon (i) by horizontally concatenating with the reference image; (iii) blending mask $M_{ij}$ with the reference image, overlaying the mask onto the reference image; and (iv) building upon (iii) by horizontally concatenating with the reference image. Overall, we found that the fourth solution (i.e., the prompt image in Fig. 3) leads to more accurate object recognition results.
Text prompt template. Our prompt template is shown in Fig. A1. Given the name of the object (e.g., chair), a specific prompt text can be derived from the carefully designed template by replacing [object_name] in the template with the name of the object. Note that all material parts share the same prompt text, rather than deriving a prompt for every material part.
Figure A2: Qualitative comparison to Text2Mesh and Paint-it. It can be seen that our approach can produce more realistic and consistent textures.
Appendix B Experimental Details
B.1 Datasets
We evaluated our method on three datasets including 3D-FUTURE [30], 3DCoMPaT [31] and Objaverse [33] dataset.
3D-FUTURE [30] is a richly-annotated large-scale dataset of 3D furniture shapes in household scenarios. It contains 9,992 unique industrial 3D CAD meshes of furniture with related attributes such as category, style, theme, etc., and high-resolution informative textures developed by professional designers.
3DCoMPaT [31] is a large-scale 3D dataset of more than 7.19 million rendered compositions of materials on parts of 7,262 unique 3D models. 3DCoMPaT covers 43 shape categories, 235 unique part names, and 167 unique material classes that can be applied to parts of 3D objects. This dataset primarily focuses on stylizing 3D shapes at part-level with compatible materials. Therefore, each object in this dataset has material part segmentation annotation.
Objaverse 1.0 [33] is a large-scale corpus of high quality, richly annotated 3D objects, which contains over 800,000 3D assets designed by over 100,000 artists. Assets in Objaverse not only belong to varied categories like animals, humans, and vehicles, but also include interiors and exteriors of large spaces that can be used, e.g., to train embodied agents. Objaverse improves upon present day 3D repositories in terms of scale, number of categories, and in the visual diversity of instances within a category.
B.2 Relighting in Blender
In the relighting experiments of the main paper (i.e., Sec. 4.3), we utilized the HDRI environment maps (copyright-free) from Poly Haven [44], a curated public asset library for visual effects artists and game designers. For the texture maps applied to the meshes in Blender, we extracted diffuse maps, normal maps, and roughness maps from the computational graphs of the optimized procedural materials, as all computational graphs can output at least these three types of maps. Moreover, for materials classified as metal, we also employed the metallic maps. Subsequently, we create a Blender Principled BSDF for each material part. The Principled BSDF in Blender is based on the OpenPBR surface shading model, and provides parameters compatible with similar PBR shaders found in other software, such as the Disney and Standard Surface models.
Appendix C Extended Experiments
In this section, we additionally compare our method with two other state of the art text-driven texture generation methods: Text2Mesh [14] and Paint-it [45]. Unlike diffusion-based methods, Text2Mesh directly optimizes textures and geometries via a CLIP-based optimization function. In addition, the mesh texture generated by Text2Mesh is saved as per-vertex color, rather than texture maps. Similar to Text2Tex [1] and TEXTure [2], we removed the geometry optimization and only optimized the textures for Text2Mesh. Paint-it [45] utilizes a neural network to represent the PBR texture maps of a mesh, and optimizes this specific neural network through a diffusion-based loss and differentiable rendering. For the texture maps generated by Paint-it [45], we constructed the PBR material nodes claimed by NVDiffRast [28] (i.e., the differentiable renderer used by Paint-it) in Blender to create the material, and then obtained the rendering results. As shown in Fig. A2, our method can produce more realistic and consistent textures. The texturing results generated by Text2Mesh lack texture information and show repetition because it relies solely on CLIP guidance to optimize per-vertex colors, as also demonstrated in Text2Tex [1] and TEXTure [2]. Paint-it utilizes a randomly initialized U-Net network as a proxy for PBR texture maps and lacks material priors, leading to reduced realism in actual application scenarios (e.g., in Blender).
Declarations
Availability of data and materials
The procedural materials utilized in this study can be downloaded from the official website of DiffMat [5, 6] at https://github.com/mit-gfx/diffmat-legacy.
Author contributions
Ziqiang Dang and Zhaopeng Cui conceived of the presented idea, Ziqiang Dang proposed and implemented the prototype system. Ziqiang Dang, Wenqi Dong and Zesong Yang designed and performed the experiments. Ziqiang Dang and Zhaopeng Cui wrote the paper. Bangbang Yang and Liang Li contributed to the method design. Yuewen Ma and Zhaopeng Cui supervised the project. Ziqiang Dang created the demo video. All authors discussed the results and contributed to the final manuscript.
Acknowledgements
This research was partially supported by the National Natural Science Foundation of China (No. 62441222), and the Information Technology Center and State Key Lab of CAD&CG, Zhejiang University. We also express our gratitude to the anonymous reviewers for their professional and constructive comments.
Declaration of competing interests
The authors have no competing interests to declare that are relevant to the content of this article.
Electronic Supplementary Material
A supplementary video containing method explanations and qualitative experimental results is available in the online version of this article.